• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Seonho Choi

    Log entry time 17:20:46 on December14,2007

    Entry number 220825

    This entry is a followup to: 220796


    keyword=Compton DAQ/computer failure

    The beam came back at around 4:25 (MCC was trying to solve Hall-C beam delivery problem.) At that time, we found out that the Compton DAQ was not working. Several tries to reboot coda failed and called A. Camsonne. He tried to reboot the machine remotely, but now the computer itself is stuck during boot process with a message "Kernel panic: VFS Unable to mount root fs ...". We tried to reset the computer using the "reset" button or even powered it off completely and powered it on, but still the computer is stuck at the same message during booting. A. Camsonne is in and working on the problem. From the halog entry 220796, it seems that the Compton DAQ was working until 13:10 this afternoon and then it stopped working suddenly while beam was on.